OpenStudy (anonymous):

simplify to the lowest terms √162 + 6√2 -√50

OpenStudy (anonymous):

Exact form: 10 sqrt 2 Decimal form: 14.142

OpenStudy (anonymous):

In this case, use the lowest (x)^(1/2), which is 2^(1/2) (81*2)^(1/2) +6(2)^(1/2)-(25*2)^(1/2) =9(2)^(1/2)+6(2)^(1/2)-5(2)^(1/2) =10(2)^1/2

OpenStudy (anonymous):

I have the same answer as well.